use two different methods to find where the following two lines intersect
y=x+8, y=-3x+4

Answer by mananth(16949)

y=x+8,
y=-3x+4
elimination method
subtract the equations
y-y= x+8-(-3x+4)
0=x+8+3x-4
0=4x+4
add -4
-4=4x+4-4
-4=4x
divide by 4
x=-1
..
plug value of x in first equation
y=-1+8
y=7
...
substitution method
y=x+8,
y=-3x+4
plug value of y in second equation
x+8=-3x+4
reorder the equation
4x=-4
x=-1
..
plug value of x in the equation
y=-1+8
y=7